Recent Log Entries

The following cavity histories have been updated: 1L12-6

Log Number: 3813797Logbook(s): Tags: Body: 

The following cavity histories have been updated:
1L12-6
Username: tdavis
Date/Time: 07-09-20_10:47:41
RF Status: Good
Tuner Status: Bad
Comment: Verified that the tuner motors are trying to turn but the tuners inside of the cryomodule are frozen. Frank Humphries also was there to verify.
Special Instructions:

The following cavity histories have been updated: 1L12-2

Log Number: 3813796Logbook(s): Tags: Body: 

The following cavity histories have been updated:
1L12-2
Username: tdavis
Date/Time: 07-09-20_10:47:41
RF Status: Good
Tuner Status: Bad
Comment: Verified that the tuner motors are trying to turn but the tuners inside of the cryomodule are frozen. Frank Humphries also was there to verify.
Special Instructions:

The following cavity histories have been updated: 1L12-1

Log Number: 3813795Logbook(s): Tags: Body: 

The following cavity histories have been updated:
1L12-1
Username: tdavis
Date/Time: 07-09-20_10:47:41
RF Status: Good
Tuner Status: Bad
Comment: Verified that the tuner motors are trying to turn but the tuners inside of the cryomodule are frozen. Frank Humphries also was there to verify.
Special Instructions:

iocsl25rfhpa restored

Log Number: 3813793Logbook(s): Tags: Body: restorelib 2.1 Target restore date was 2020-07-09 10:31:51 The output from the myRestore execution was as follows: iocsl25rfhpa Rf12Hpa@2020-07-09 10:31 111(111) restored system@2020-07-09 10:31 2(2) restored Execute: /usr/csite/certified/bin/ced2Epics -ced ced -A -s -i iocsl25rfhpa (success) Procedure return status: Success

2L03 vacuum equipment

Log Number: 3813792Logbook(s): Body: 

Beamline pump isolated & removed from VIP2L03B right angle valve. Cleanroom and all equipment
removed from the area.

Follow-up Re: qs Utility woes

Log Number: 3813788Logbook(s): Body: 

The harp may have hung up. When there is an opportunity, we should try to manually operate the harp. Perhaps a reboot of iocin1 is in fact necessary. There are quite a bit of Injector devices attached to iocin1, so the reboot should not be done casually.

PD 8am Meeting Summary

Log Number: 3813770Logbook(s): Body: 

HCO downstairs work through noon:
- 1L12 including tuners 1 & 6 - multiple groups coordinating to troubleshoot
- ARC9 shunt modules check

1L25 Decarad setup will stay in place until long SAD

Rad levels elevated at end of NL - RWP in works

Tunnel work may be crowded at times (especially around 1L12, 1L25 and 2L03). Maintain COVID distancing at all times.

1L10 cavities 2 and 3

Log Number: 3813768Logbook(s): Entry Makers: mountsBody: 

Cav 2 has no GMES - everything from cables to RFCM checked. Cavity 3 will be looked at today...

Pages